OpenStudy (anonymous):

Given the function f(x) = x+8 find the following f(6)=, f(-6) = and f(0) =

OpenStudy (anonymous):

6=x+8, -6=x+8 ,0=x+8

OpenStudy (anonymous):

just solve for x